# SynFlowNet: Design of Diverse and Novel Molecules with Synthesis Constraints

### v2

@@ Abstract
Generative models see increasing use in computer-aided drug design. However, while performing well at capturing distributions of molecular motifs, they often produce synthetically inaccessible molecules. To address this, we introduce SynFlowNet, a GFlowNet model whose action space uses chemical reactions and buyable reactants to sequentially build new molecules. By incorporating forward synthesis as an explicit constraint of the generative mechanism, we aim at bridging the gap between in silico molecular generation and real world synthesis capabilities. We evaluate our approach using synthetic accessibility scores and an independent retrosynthesis tool to assess the synthesizability of our compounds, and motivate the choice of GFlowNets through considerable improvement in sample diversity compared to baselines. Additionally, we identify challenges with reaction encodings that can complicate traversal of the MDP in the backward direction. To address this, we introduce various strategies for learning the GFlowNet backward policy and thus demonstrate how additional constraints can be integrated into the GFlowNet MDP framework. This approach enables our model to successfully identify synthesis pathways for previously unseen molecules.
Source code available at https://github.com/mirunacrt/synflownet
@@ 1 Introduction
Figure 1: SynFlowNet allows for synthesis-aware molecule generation. (A) The state space is induced by combining purchasable building blocks (BBs) and chemical reactions. Every final molecule (rectangle box) is associated with a reward. Training trajectories are constructed by sampling the model forward. Our policy $P_{F}(a|s)$ is parameterized as a graph transformer which at each timestep processes the current molecular state $s_{t}$ and outputs a shared embedding which is passed to separate MLP heads to predict the action logits for different action types (5 forward and 3 backward action types). An action $a_{t}$ is then sampled from this hierarchical distribution to transition to the next state via a reaction. Newly added BBs are highlighted in yellow. (B) To allow handling large sets of BBs (up to 200k), we represent them using Morgan fingerprints and compute the probability of sampling a particular BB from the normalised dot product between this representation and the embedding vector of the current state. (C) Finally, when traversing the MDP backwards, to reduce the probability of exiting the MDP defined by our set of reactions and BBs, we train the backward policy to avoid paths that do not terminate in the initial state $s_{0}$.
@@
  • We identify an inherent problem of employing a reaction-based Markov Decision Process (MDP) with GFlowNets, stemming from the lack of guarantee that backward-constructed trajectories can return to $s_{0}$. To resolve this issue, we propose one of the first attempts at training the backward policy in a GFlowNet with a separate objective from for the forward policy, which we show to correct backward-generated trajectories and improve sample quality and diversity.
@@ 2.1 GFlowNets
Bengio et al. (2021) have used GFlowNets to generate molecules with high binding affinity to a protein target by linking fragments to form a junction tree (Jin et al., 2019). For generative chemistry, the framework was extended to multi-objective optimisation (Jain et al., 2023; Roy et al., 2023), where the model was trained to simultaneously optimise for binding affinity to the protein target, Synthetic Accessibility (SA), drug likeness (QED) and molecular weight. Recently, Shen et al. (2023) extended the framework to pocket-conditioned molecular generation, employing the same action space to generate molecules conditioned on different protein targets.
2.2 GFlowNets with a trained backward policy
@@ 2.3 Generative models for molecule design
A large number of works have been proposed for generative molecular design (Du et al., 2024). Early methods employed techniques such as variational autoencoders (VAEs) (Gómez-Bombarelli et al., 2018), deep reinforcement learning (Segler et al., 2017; Olivecrona et al., 2017b), and generative adversarial networks (GANs) (Guimaraes et al., 2018; Cao & Kipf, 2022). More recently, methods like normalizing flows (Zang & Wang, 2020) and diffusion models (Hoogeboom et al., 2022; Schneuing et al., 2023; Harris et al., 2023b; Torge et al., 2023; Igashov et al., 2024) have gained significant attention due to their ability to model complex distributions more effectively. Despite these advances, challenges remain, particularly in ensuring that generated molecules adhere to physical and chemical constraints (Stanley & Segler, 2023; Harris et al., 2023a).
@@ 2.4 Synthesis-aware molecule generation
The idea of tackling molecule generation and synthesis simultaneously has been investigated early by Vinkers et al. (2003), who introduced SYNOPSIS, which generates molecules from a starting dataset of available compounds, relying on applying chemical modifications to functional groups and assessing the value of the product with a fitness function. The works of Bradshaw et al. (2019) and Korovina et al. (2020) provided early neural models for one-step synthetic pathways. Bradshaw et al. (2020) generalized this idea as a generative model for synthesis DAGs, which can be optimized in a VAE or RL setup. Gottipati et al. (2020) used reinforcement learning to generate compounds from reactions and commercially available reactants. Gao et al. (2021) formulate an MDP to model the generation of synthesis trees, which can be optimized with respect to the desired properties of a product molecule. Luo et al. (2024) proposed a model that can project unsynthesizable molecules from existing generative models to synthesizable chemical space by utilizing postfix notations to represent synthesis pathways. Guo & Schwaller (2024) showed that retrosynthesis models can be treated as an oracle in goal-directed molecule generation. Concurrently to our work, Koziarski et al. (2024) propose a similar GFlowNet-based framework for synthesizable molecular generation but using a different set of reaction templates and making use of a different backward policy.
While our work resembles an RL setting for synthesis-aware molecular generation (Gottipati et al., 2020; Horwood & Noutahi, 2020), the key difference lies in the sampling distribution of the learned model. Contrary to RL, the GFlowNet objective is not to generate the single highest-return sequence of actions, but rather to maximise both performance and diversity by sampling terminal states proportionally to their reward. This is especially useful in the context of molecule generation, where we want to explore different modes of the distribution of interest.
@@ 3.1 MDP of chemically accessible space
Data
We use commercially available building blocks (BBs) from Enamine , which are small fragments of molecules prepared in bulk to be readily synthesised into candidate molecules. Reaction templates are obtained from two publicly available template libraries (Button et al., 2019; Hartenfeller et al., 2012). After preprocessing, we obtain a total of 105 templates: 13 uni- and 92 bi-molecular reactions respectively (see App. A.2). We also use 12 Enamine REAL reactions in a small number of experiments (see Section 4.6 and App. A.2).
@@ 3.2 Challenges with Generating Backward Paths
In that setting, we (1) generate trajectories using $P_{F}$, (2) update $P_{F}$ according to the trajectory balance objective in Equation 1 and (3) update $P_{B}$ using these same trajectories according to Equation 2 (see see Algorithm 1 in Appendix A.4).
@@
Interestingly, training the backward policy to navigate back to $s_{0}$ is analogous to the retrosynthesis problem (Corey & Cheng, 1989; Segler et al., 2018). We employ this approach of training $P_{B}$ against a different objective than $P_{F}$ to palliate to a design challenge of the MDP, but a similar strategy could also be employed to fold additional preferences over different synthesis routes leading to the same terminal state into the system, for example to take into account the synthesis costs of a particular path. This possibility is left to future work.
@@ 3.3 Scaling of the reactant space to large number of building blocks
To ensure synthetic accessibility of all samples, our model is inherently constrained by the initial set of available building blocks (BBs). To cover a large chemical space, it is thus crucial to use an extensive BB collection. The model must demonstrate scalability to accommodate larger sets of BBs, both in terms of training efficiency and overall performance. To do so, we change the representation of the BBs and their selection mechanism, as shown in Figure 1B. Instead of the weight matrix of the mapping from hidden units to logits associated with BBs to be randomly initialized, it is fixed to be the matrix of binary Morgan fingerprints (Rogers & Hahn, 2010).
@@ 4.1 Reaction-based MDP design
Figure 3: Example of synthesis pathway generated by SynFlowNet. The reward here is the sEH binding affinity proxy. Right shows the Vina docked pose and molecule metrics.
Figure 4: Comparison across MDPs. We evaluate four GFlowNet models: SynFlowNet is trained with an action space of chemical reactions and maximum trajectory lengths (L) of 3 and 4, FragGFN and FragGFN SA are trained with an action space of fragments, however the latter also optimises for synthetic accessibility (SA), besides the sEH binding proxy reward. SynFlowNet molecules are achieving higher binding scores and better synthesisability.
@@
The results in Fig. 4 show that our MDP design achieves great improvement in terms of synthesisability, while preserving high rewards and diversity. We see that within the reaction environment, MDP design choices can further influence sample quality: a maximum trajectory length of 3 achieves better SA scores and AiZynthFinder (Genheden et al., 2020) retrosynthesis success (62%) compared to a maximum trajectory length of 4 (40%) [^1] The reactions available to AiZynthFinder are the US patent office (USPTO) set (Lowe, ), which are different to the reactants used by SynFlowNet, so one should not expect a 100% success score.. Note that synthesisability metrics are correlated with molecule size (Skoraczyński et al., 2023), and that molecules assembled from longer synthesis routes are naturally larger (see Table A.2). Both FragGFN models score 0% with AiZynthFinder. Diversity, measured as average Tanimoto distances between molecular fingerprints, is preserved from a fragments to a reaction environment, albeit less so for shorter trajectories. Figure 3 is an example of output from SynFlowNet, returning also the synthesis pathway extracted from the MDP.
@@ 4.2 GFlowNets as samplers of chemical space
In Fig. 5, we show the reward distribution of generated samples after equal numbers of training steps. We notice that soft Q-learning collapses to sampling a narrower distribution of high-reward molecules for all targets. This is also shown in the number of Bemis-Murcko scaffolds counted for molecules with rewards above a certain threshold, where soft Q-learning quickly collapses (note the logarithmic scale of the y-axis).
@@ 4.3 Further comparison to baselines
We continue by comparing SynFlowNet to strong baselines from the literature. First, REINVENT uses a policy-gradient method to tune an RNN pre-trained to generate SMILES strings (Blaschke et al., 2020; Loeffler et al., 2024). REINVENT has been shown to outperform many models in terms of sample efficiency (Gao et al., 2022) and proposing realistic 3D molecules upon docking (Ciepliński et al., 2023). Second, SyntheMol shares a similar MDP design with SynFlowNet, but is a search-based method using Monte Carlo Tree Search (Swanson et al., 2024). Contrary to our model, synthesis pathways in SyntheMol contain one reaction, and its estimated state space has 30 billion molecules, reached from the Enamine building block set and 13 reactions (Swanson et al., 2024).
The comparison is summarized in Figure 6. SynFlowNet achieves comparable rewards to REINVENT, and better sample diversity. Although REINVENT does not use explicit knowledge of synthesizability in its generation process, it implicitly optimizes for it by being a likelihood model overfit on a curated ChEMBL dataset (Zdrazil et al., 2023a). This is reflected in our last sub-plot, where we look at the maximum similarity to ChEMBL molecules, as a proxy for novelty. This proxy is strong enough, as ChEMBL contains one of the largest collection of diverse drug-like molecules to date (Zdrazil et al., 2023b). While REINVENT seems to be competitive in terms of synthesizability and reward, it generates molecules with poor novelty, staying close to its pretraining distribution. Notably, SynFlowNet achieves high novelty for high-scoring molecules. Overall, SyntheMol scores comparably to SynFlowNet, however SynFlowNet achieves a better reward/diversity trade-off and SynFlowNet’s capabilities are extended to explore larger state spaces (see Fig. 2). We repeat the comparison on other targets, and again find that SynFlowNet outperforms SyntheMol (see App. B).
@@ 4.4 Improved MDP consistency through trained backward policy
We evaluate the effectiveness of our proposed training schemes for the backward policy $P_{B}$ by measuring (1) whether $P_{B}$ can ensure that backward-constructed trajectories reliably find a trajectory back to the initial state $s_{0}$ and (2) whether it brings any benefit to the forward policy. In Table 1 we compare a fixed, uniform backward policy to three versions of a parameterized backward policy: a free policy, updated w.r.t. the trajectory balance loss, (see Eq. 1), a policy trained with maximum likelihood on the forward-generated trajectories and a policy that is allowed to explore the backward action space and trained with REINFORCE to find paths leading back to $s_{0}$. We observe that both the maximum likelihood and REINFORCE policies succeed in ensuring that backward flow is not lost outside of the MDP, as they manage to construct trajectories that start from terminal states sampled from $P_{F}$ all the way to $s_{0}$. We refer to such routes as solved routes (train) in Table 1, as the terminal states have been visited by the GFlowNet during training. We also test the ability of the trained policies to retrieve synthesis routes for molecules which have not been visited during training ($test$). These were obtained from a random sampler using our set of reactions and building blocks. While we see small differences in the number of high-reward modes discovered for different backward policies, we see that a free policy consistently fails to be competitive with the rest. Overall, the maximum likelihood and REINFORCE policies prove effective in guiding $P_{F}$ to high reward modes, while ensuring MDP consistency.
@@
Table 1: Effect of different training paradigms for $\boldsymbol{P_{B}}$. Training the GFN ${P_{B}}$ ensures that backward-constructed trajectories belong to our MDP and can marginally improve ${P_{F}}$. % of solved routes refers to the ability of $P_{B}$ to retrieve synthesis routes for on-policy (train) and off-policy (test) molecules reachable through our state space. Test molecules have not been seen during training. # of high reward ($R>0.9$) modes from $P_{F}$ is reported out of $\sim 500,000$ samples seen during training.

### v3

@@ Abstract
Generative models see increasing use in computer-aided drug design. However, while performing well at capturing distributions of molecular motifs, they often produce synthetically inaccessible molecules. To address this, we introduce SynFlowNet, a GFlowNet model whose action space uses chemical reactions and purchasable reactants to sequentially build new molecules. By incorporating forward synthesis as an explicit constraint of the generative mechanism, we aim at bridging the gap between in silico molecular generation and real world synthesis capabilities. We evaluate our approach using synthetic accessibility scores and an independent retrosynthesis tool to assess the synthesizability of our compounds, and motivate the choice of GFlowNets through considerable improvement in sample diversity compared to baselines. Additionally, we identify challenges with reaction encodings that can complicate traversal of the MDP in the backward direction. To address this, we introduce various strategies for learning the GFlowNet backward policy and thus demonstrate how additional constraints can be integrated into the GFlowNet MDP framework. This approach enables our model to successfully identify synthesis pathways for previously unseen molecules. Source code is available at https://github.com/mirunacrt/synflownet.
@@ 1 Introduction
Figure 1: SynFlowNet allows for synthesis-aware molecule generation. (A) The state space is induced by combining purchasable building blocks (BBs) and chemical reactions. Every final molecule (rectangle box) is associated with a reward. Training trajectories are constructed by sampling the model forward. Our policy $P_{F}(a|s)$ is parameterized as a graph transformer which at each timestep processes the current molecular state $s_{t}$ and outputs a shared embedding which is passed to separate MLP heads to predict the action logits for different action types (5 forward and 3 backward action types). An action $a_{t}$ is then sampled from this hierarchical distribution to transition to the next state via a reaction. (B) To allow handling large sets of BBs (up to 200k), we represent them using Morgan fingerprints and compute the probability of sampling a particular BB from the normalised dot product between this representation and the MLP output for the current state. The state representation is concatenated with the one-hot encoding of the selected reaction. (C) Finally, when traversing the MDP backwards, to reduce the probability of exiting the MDP defined by our set of reactions and BBs, we train the backward policy to avoid paths that do not terminate in $s_{0}$.
@@
  • We identify an inherent problem of employing a reaction-based Markov Decision Process (MDP) with GFlowNets, stemming from the lack of guarantee that backward-constructed trajectories can return to the initial state $s_{0}$. To resolve this issue, we propose one of the first attempts at training the backward policy in a GFlowNet with a separate objective from the forward policy, which we show to correct backward-generated trajectories in the MDP while retaining the ability to discover diverse and high-reward modes.
@@ 2.1 GFlowNets
Bengio et al. (2021) have used GFlowNets to generate molecules with high binding affinity to a protein target by linking fragments to form a junction tree (Jin et al., 2019). For generative chemistry, the framework was extended to multi-objective optimisation (Jain et al., 2023; Roy et al., 2023), where the model was trained to simultaneously optimise for binding affinity to the protein target, Synthetic Accessibility (SA), drug likeness (QED) and molecular weight.
2.2 GFlowNets with a parameterized backward policy
@@ 2.3 Generative models for molecule design
A large number of works have been proposed for generative molecular design (Du et al., 2024). Early methods employed techniques such as variational autoencoders (VAEs) (Gómez-Bombarelli et al., 2018), deep reinforcement learning (Segler et al., 2017; Olivecrona et al., 2017b), and generative adversarial networks (GANs) (Guimaraes et al., 2018; Cao & Kipf, 2022). (Zang & Wang, 2020) Despite these advances, challenges remain, particularly in ensuring that generated molecules adhere to physical and chemical constraints (Stanley & Segler, 2023; Harris et al., 2023).
@@ 2.4 Synthesis-aware molecule generation
The idea of tackling molecule generation and synthesis simultaneously has been investigated early by Vinkers et al. (2003), who introduced SYNOPSIS, which generates molecules from a starting dataset of available compounds, relying on applying chemical modifications to functional groups and assessing the value of the product with a fitness function. The works of Bradshaw et al. (2019) and Korovina et al. (2020) provided early neural models for one-step synthetic pathways. Bradshaw et al. (2020) generalized this idea as a generative model for synthesis DAGs, which can be optimized in a VAE or RL setup. Gottipati et al. (2020b) used reinforcement learning to generate compounds from reactions and commercially available reactants. Gao et al. (2021) formulate an MDP to model the generation of synthesis trees, which can be optimized with respect to the desired properties of a product molecule. Luo et al. (2024) proposed a model that can project unsynthesizable molecules from existing generative models to synthesizable chemical space by utilizing postfix notations to represent synthesis pathways. Guo & Schwaller (2024) showed that retrosynthesis models can be treated as an oracle in goal-directed molecule generation. Concurrently to our work, Koziarski et al. (2024) propose a similar GFlowNet-based framework for synthesizable molecular generation but using a different set of reaction templates and making use of a different backward policy.
While our work resembles an RL setting for synthesis-aware molecular generation (Gottipati et al., 2020b; Horwood & Noutahi, 2020), the key difference lies in the sampling distribution of the learned model. Contrary to RL, the GFlowNet objective is not to generate the single highest-return sequence of actions, but rather to maximise both performance and diversity by sampling terminal states proportionally to their reward. This is especially useful in the context of molecule generation, where we want to explore different modes of the distribution of interest.
@@ 3.2 Challenges with Generating Backward Paths
In that setting, we (1) generate trajectories using $P_{F}$, (2) update $P_{F}$ according to the trajectory balance objective in Equation 1 and (3) update $P_{B}$ using these same trajectories according to Equation 2 (see Algorithm 1 in Appendix A.4).
@@
Interestingly, training the backward policy to navigate back to $s_{0}$ is analogous to the retrosynthesis problem (Corey & Cheng, 1989; Segler et al., 2018). We employ this approach of training $P_{B}$ against a different objective than $P_{F}$ to palliate to a design challenge of the MDP, but a similar strategy could also be employed to fold additional preferences over different synthesis routes leading to the same terminal state into the system, for example to take into account the synthesis costs of a particular path.
@@ 3.3 Scaling of the reactant space to large number of building blocks
To ensure synthetic accessibility of all samples, our model is inherently constrained by the initial set of available building blocks (BBs). To cover a large chemical space, it is thus crucial to use an extensive BB collection. The model must demonstrate scalability to accommodate larger sets of BBs, both in terms of training efficiency and overall performance. To do so, we follow Dulac-Arnold et al. (2015); Gottipati et al. (2020a) and change the representation of the BBs and their selection mechanism, as shown in Figure 1B. Instead of the weight matrix of the mapping from hidden units to logits associated with BBs to be randomly initialized, it is fixed to be the matrix of binary Morgan fingerprints (Rogers & Hahn, 2010).
@@ 4.1 Reaction-based MDP design
Figure 3: Comparison across MDPs. We evaluate four GFlowNet models: SynFlowNet is trained with an action space of chemical reactions and maximum trajectory lengths (L) of 3 and 4, FragGFN and FragGFN SA are trained with an action space of fragments, however the latter also optimises for synthetic accessibility (SA), besides the sEH binding proxy reward. SynFlowNet molecules are achieving higher binding scores and better synthesizability.
@@
The results in Fig. 3 show that our MDP design achieves great improvement in terms of synthesizability, while preserving high rewards and diversity. While one might expect the more expressive FragGFN model to result in higher rewards, we hypothesize that the larger exploration space in the fragments environment (see Figure 2) hinders the model’s efficiency. We see that within the reaction environment, MDP design choices can further influence sample quality: a maximum trajectory length of 3 achieves better SA scores and AiZynthFinder (Genheden et al., 2020) retrosynthesis success (62%) compared to a maximum trajectory length of 4 (40%). Note that synthesisability metrics are correlated with molecule size (Skoraczyński et al., 2023), and that molecules assembled from longer synthesis routes are naturally larger (see Table A.3). Both FragGFN models score 0% with AiZynthFinder. Diversity, measured as average Tanimoto distances between molecular fingerprints, is preserved from a fragments to a reaction environment, albeit less so for shorter trajectories.
@@ 4.2 GFlowNets as samplers of chemical space
In Fig. 4, we show the reward distribution of generated samples after equal numbers of training steps. We notice that soft Q-learning collapses to sampling a narrower distribution of high-reward molecules for all targets. This is also shown in the number of Bemis-Murcko scaffolds counted for molecules with rewards above a certain threshold, where soft Q-learning quickly collapses.
@@ 4.3 Further comparison to baselines
We continue by comparing SynFlowNet to strong baselines from the literature. First, REINVENT uses a policy-gradient method to tune an RNN pre-trained to generate SMILES strings (Blaschke et al., 2020; Loeffler et al., 2024). REINVENT has been shown to outperform many models in terms of sample efficiency (Gao et al., 2022a) and proposing realistic 3D molecules upon docking (Ciepliński et al., 2023). Second, SyntheMol shares a similar MDP design with SynFlowNet, but is a search-based method using Monte Carlo Tree Search (Swanson et al., 2024). Contrary to our model, synthesis pathways in SyntheMol contain one reaction, and its estimated state space has 30 billion molecules, reached from the Enamine building block set and 13 reactions (Swanson et al., 2024).
The comparison is summarized in Figure 5. SynFlowNet achieves comparable rewards to REINVENT, and better sample diversity. Although REINVENT does not use explicit knowledge of synthesizability in its generation process, it implicitly optimizes for it by being a likelihood model overfit on a curated ChEMBL dataset (Zdrazil et al., 2023). This is reflected in our last sub-plot, where we look at the maximum similarity to ChEMBL molecules, as a proxy for novelty. This proxy is strong enough, as ChEMBL contains one of the largest collection of diverse drug-like molecules to date (Zdrazil et al., 2023). While REINVENT seems to be competitive in terms of synthesizability and reward, it generates molecules with poor novelty, staying close to its pretraining distribution. Notably, SynFlowNet achieves high novelty for high-scoring molecules, even when trained using ChEMBL-derived building blocks (see App. Fig. A.5). Overall, SyntheMol scores comparably to SynFlowNet, however SynFlowNet achieves a better reward/diversity trade-off and SynFlowNet’s capabilities are extended to explore larger state spaces (see Fig. 2). We repeat the comparison on other targets, and again find that SynFlowNet outperforms SyntheMol (see App. B.1). Further experiments showcasing SynFlowNet’s sample efficiency and performance when optimizing for direct docking score calculations, are discussed in Appendix B.1. We again discover that SynFlowNet is competitive to other models and that it improves sample efficiency over the fragments GFlowNet.
@@
4.3.1 Rediscovery tasks
We investigate whether SynFlowNet can discover known synthesizable actives by training it with rediscovery rewards (Huang et al., 2021). The results are shown in Table LABEL:tab:rediscovery, where we report top-k similarity and SA scores. When running the SpaceLight software (BioSolveIt, 2024) to determine whether the two molecules are present in the REAL space (i.e. reachable through our set of building blocks), we found that aripiprazole is present and that celecoxib is not, which explains the two rates.
Table 1: SynFlowNet performance on rediscovery tasks. We report mean and standard deviation of top-10 discovered molecules from running three models with different seeds.
Task Reward ($\uparrow$) SA ($\downarrow$)
Aripiprazole rediscovery $0.90\pm 0.00$ $2.19\pm 0.00$
Celecoxib rediscovery $0.48\pm 0.01$ $2.47\pm 0.04$
@@ 4.4 Improved MDP consistency through trained backward policy
We evaluate the effectiveness of our proposed training schemes for the backward policy $P_{B}$ by measuring (1) whether $P_{B}$ can ensure that backward-constructed trajectories reliably find a trajectory back to the initial state $s_{0}$ and (2) whether it brings any benefit to the forward policy. In Table 2 we compare a fixed, uniform backward policy to three versions of a parameterized backward policy: a free policy, updated w.r.t. the trajectory balance loss, (see Eq. 1), a policy trained with maximum likelihood on the forward-generated trajectories and a policy that is allowed to explore the backward action space and trained with REINFORCE to find paths leading back to $s_{0}$. We observe that both the maximum likelihood and REINFORCE policies succeed in ensuring that backward flow is not lost outside of the MDP, as they manage to construct trajectories that start from terminal states sampled from $P_{F}$ all the way to $s_{0}$. We refer to such routes as solved routes (train) in Table 2, as the terminal states have been visited by the GFlowNet during training. We also test the ability of the trained policies to retrieve synthesis routes for molecules which have not been visited during training ($test$). These were obtained from a random sampler using our set of reactions and building blocks. While we see small differences in the number of high-reward modes discovered for different backward policies, we see that a free policy consistently fails to be competitive with the rest. More ablation studies in App. A.6 show that an entropy-regularized REINFORCE policy excels at discovering high reward modes. Overall, the maximum likelihood and REINFORCE policies prove effective in guiding $P_{F}$ to high reward modes, while ensuring MDP consistency, which is crucial for sampling from the Boltzmann distribution defined by the reward function $R$ (Bengio et al., 2023).
@@
Table 2: Effect of different training paradigms for $\bm{P_{B}}$. Training the GFN ${P_{B}}$ ensures that backward-constructed trajectories belong to our MDP and can marginally improve ${P_{F}}$. % of solved routes refers to the ability of $P_{B}$ to retrieve synthesis routes for on-policy (train) and off-policy (test) molecules reachable through our state space. Test molecules have not been seen during training. # of high reward ($R>0.9$) modes from $P_{F}$ is reported out of $\sim 500,000$ samples seen during training.